Your goal is to be able to withdraw $4,100 for each of the next nine years beginning one year from today. The return on the investment is expected to be 11%. The amount that needs to be invested today is closest to: (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use appropriate factor(s) from the tables provided.)

$36,900.

$39,096.

$30,742.

$22,702
Answer #1

Factor using annuity formula =1*((1-(1+11%)^-9)/11%) =5.53705
PV of amount invested =PMT*PV of factor =4100*5.53705 =22701.91 or 22702